# A NEW POLYNOMIAL FOR CHECKERBOARD-COLORABLE 4-VALENT VIRTUAL GRAPHS

### v1

@@ Abstract
We assign a new polynomial to any checkerboard-colorable 4-valent virtual graph in terms of its Euler circuit expansion. This provides a new combinatorial formulation of the Kauffman-Jones polynomial for checkerboard-colorable virtual links.
footnotetext: 2020 Mathematics Subject Classification. 57K10, 57K14, 57K18 footnotetext: Key words and phrases. 4-valent graphs, Jones-Kauffman polynomial, Euler circuits, virtual links, checkerboard-colorable.
@@ 2 Preliminary
A 4-valent graph has a source target structure iff it is checkerboard-colorable. In fact, if $G$ is a 4-valent graph with a source target structure, then fix one of the two possible orientations of the edges of $G$ induced by the source target structure. To each oriented edge, we associate a normal vector satisfying the right hand rule in the plane as depicted in the left of Figure 3. By coloring the side of each edge containing its normal vector we get a checkerboard coloring of $G$. Conversely, If $G$ is a checkerboard colored 4-valent graph, then, for each edge, we put a normal vector inside its colored side with initial point on that edge as depicted in the right of Figure 3. Then we orient the edge by using the right hand rule.
Throughout the remainder of this paper, we adopt the convention that a checkerboard coloring is associated with a source-target orientation as described above, with edges oriented according to the right-hand rule and colored based on the direction of their normal vectors.
@@
A 4-valent graph $G$ is said to be checkerboard-embedded in some orientable surface $\Sigma$ if there exists an embedding $e$ of $G$ in $\Sigma$ such that the components of $\Sigma\setminus e(G)$ are 2-cells and there exists a coloring of these components by two colors, say black and white, such that two components of $\Sigma\setminus e(G)$ which are adjacent by an edge of $e(G)$ have always distinct colors. Manturov in [14, 15] showed that a 4-valent virtual graph $G$ is checkerboard-colorable iff it is checkerboard-embeddable.
@@ 2.1 Euler circuits and chord diagrams
From now on, we assume that any $2$-digraph $G$ is checkerboard-colorable unless mentioned otherwise. In the graph $G$, we let $v_{1},\ldots,v_{n}$ be the set of vertices of such graph and $\gamma$ be an Euler circuit of $G$. Draw a simple circle $C$ in the plane and mark $2n$ evenly spaced points around its circumference. Start traveling $\gamma$ following the orientation of the edges starting from any vertex. Label each of the $2n$ points on the circle $C$ successively with the label of the vertex we encounter. Each vertex label appears twice at exactly two distinct points on the circle. For each vertex $v_{i}$ in $G$, draw a chord inside the circle $C$ connecting the two points labeled $v_{i}$. We ensure that each pair of chords intersect at most once and no more than two chords intersect at the same point. The circle $C$ together with the $n$ chords constitute the chord diagram of $\gamma$, which we denote by $C(\gamma)$. See Figure 4 as an example for an Euler circuit $\gamma$ and its corresponding chord diagram $C(\gamma)$.
@@ 2.3 The Jones-Kauffman polynomial of a virtual link
For an oriented virtual link diagram $L$, we let $x(L)$ denotes the number of negative crossings and $y(L)$ denotes the number of positive crossings in $L$ according to the scheme in Figure 7. The writhe of the virtual link diagram $L$ is defined to be the integer $w(L)=y(L)-x(L)$.
@@
$f_{L}(q)=(-q)^{-3w(L)}\left\langle L\right\rangle\in\mathbb{Z}[q,q^{-1}],$
@@
$L$                       $L_{0}$                      $L_{1}$
Figure 8: The link diagram $L$ at the crossing $c$ and its smoothings $L_{0}$ and $L_{1}$ respectively.
@@ 3 The main results and an application
We say that a $2$-digraph is vertex-signed if each vertex is equipped with a sign $+$ or --. At each vertex $v_{i}$, we denote the two incoming edges by $e_{i_{1}}$ and $e_{i_{2}}$, and we denote the two outgoing edges by $e_{i_{3}}$ and $e_{i_{4}}$. We assume that at a fixed vertex $v_{i}$ the Euler circuit $\gamma$ of $G$ traverses successively the edges $e_{i_{1}}$ and $e_{i_{3}}$ in the first visit, and then it traverses successively $e_{i_{2}}$ and $e_{i_{4}}$ in the second visit. Merge the edges $e_{i_{1}}$ and $e_{i_{3}}$, i.e., replace them with a single new edge going from the tail of $e_{i_{1}}$ to the head of $e_{i_{3}}$. Also, we merge the edges $e_{i_{2}}$ and $e_{i_{4}}$ using the same technique and we delete the vertex $v_{i}$ to get a new graph with one less vertex. In the vicinity of the deleted vertex $v_{i}$, put a marker connecting the traces of the deleted vertex on each one of the new edges. We label the marker by $A$, $a$, $B$, or $b$ according to the coloring, the sign of the vertex and the order of the edges when traversing the path as explained in Figure 9. We iterate this merging process until all the vertices are deleted and we obtain a single immersed circle in the plane with $n$ labeled markers in the vicinities of the deleted vertices. We call such data the $\gamma$-state of the corresponding Euler circuit $\gamma$ of $G$. In Figure 10, we depict the $\gamma$-state of the Euler circuit $\gamma$ given in Figure 4.
Figure 9: The “$i$-th marker” corresponding to the deleted vertex $v_{i}$ and the rule of its labeling.
@@
where $G_{1},\ldots,G_{m}$ are the connected components of the graph $G$. Now we state the main theorem of this paper whose proof will be given in the next section.
@@
We enclose this subsection by stating the following theorem whose proof will be given in the next section which implies that the polynomial $X_{G}(q)$ satisfies a skein relation. For this end, we let $G_{v}$ be a checkerboard-colorable vertex-signed $2$-digraph with a fixed vertex $v$ depicted in Figure 11. The digraphs resulting from the two possible merging operations at $v$ described in Section 3 will be denoted by $G_{0}$ and $G_{1}$ respectively according to the scheme in Figure 11.
Figure 11: The 2-digraph $G_{v}$ at the fixed vertex $v$ and the corresponding 2-digraphs $G_{0}$ and $G_{1}$ respectively obtained by the two merging operations.
@@
    If $v$ is equipped with positive sign, then
    $X_{G_{v}}(q)=qX_{G_{0}}(q)+q^{-1}X_{G_{1}}(q).$
@@
    If $v$ is equipped with negative sign, then
    $X_{G_{v}}(q)=q^{-1}X_{G_{0}}(q)+qX_{G_{1}}(q).$
As an application, we relate the polynomial introduced in the this section to the established Jones-Kauffman polynomial of checkerboard-colorable virtual links. We start by recalling a brief background on virtual knot theory focusing on checkeroboard-colorable virtual links. Given a virtual link diagram, one can obtain a 4-valent virtual graph by replacing each classical crossing with a vertex and each virtual crossing by a virtual double point. The obtained graph is called the shadow graph of the given virtual link diagram. If this graph is checkerboard-colorable, then the associated virtual link diagram is checkerboard-colorable. For instance, the virtual link diagram in Figure 5, whose shadow graph is exactly the graph depicted in Figure 1, is checkerboard-colorable. A notable fact about the checkerboard-colorability is that it behaves well under generalized Reidemeister moves since any two equivalent checkerboard-colorable virtual link diagrams are related by a finite sequence of moves which preserve the checkerboard-colorability as shown in [8, Theorem 2.9]. A virtual link is checkerboard-colorable if it admits a checkerboard-colorable diagram. We point out that the notion of checkerboard-colorable virtual links was first introduced by Kamada in [10, 11].
@@
Let $L$ be an oriented checkerboard-colorable virtual link. Let $G$ be the shadow graph of a checkerboard-colorable diagram of $L$, then $f_{L}(q)=(-q)^{-3w(L)}X_{G}(q)$, where $f_{L}(q)$ and $w(L)$ are the Jones-Kauffman polynomial and the writhe of the oriented virtual link $L$, respectively.
@@ 4 Proof of the Main Theorems
    The third equivalence follows from that fact that the edges of $H^{v_{i}v_{j}}$ connecting $v_{h}$ to $v_{k}$ are:
    • $v_{h}$ is $v_{i}$
@@
      $\begin{cases}\mu_{i}(\gamma)\cdot\mu_{i+1}(\gamma)=\mu_{i}(\gamma_{v_{i}v_{i+1}})\cdot\mu_{i+1}(\gamma_{v_{i}v_{i+1}})&\\ \mu_{i}(\gamma^{v_{i}v_{i+1}})\cdot\mu_{i+1}(\gamma^{v_{i}v_{i+1}})=\mu_{i}((\gamma^{v_{i}v_{i+1}})_{v_{i}v_{i+1}})\cdot\mu_{i+1}((\gamma^{v_{i}v_{i+1}})_{v_{i}v_{i+1}})&\end{cases}$
      $\quad\text{or}\quad\begin{cases}\mu_{i}(\gamma)\cdot\mu_{i+1}(\gamma)=\mu_{i}((\gamma^{v_{i}v_{i+1}})_{v_{i}v_{i+1}})\cdot\mu_{i+1}((\gamma^{v_{i}v_{i+1}})_{v_{i}v_{i+1}})&\\ \mu_{i}(\gamma^{v_{i}v_{i+1}})\cdot\mu_{i+1}(\gamma^{v_{i}v_{i+1}})=\mu_{i}(\gamma_{v_{i}v_{i+1}})\cdot\mu_{i+1}(\gamma_{v_{i}v_{i+1}})&\end{cases}$
@@
      $\begin{cases}\mu_{i}(\gamma)\cdot\mu_{i+1}(\gamma)=-\mu_{i}(\gamma^{v_{i}v_{i+1}})\cdot\mu_{i+1}(\gamma^{v_{i}v_{i+1}})&\\ \mu_{i}(\gamma_{v_{i}v_{i+1}})\cdot\mu_{i+1}(\gamma_{v_{i}v_{i+1}})=-\mu_{i}((\gamma^{v_{i}v_{i+1}})_{v_{i}v_{i+1}})\cdot\mu_{i+1}((\gamma^{v_{i}v_{i+1}})_{v_{i}v_{i+1}})&\end{cases}$
@@
      $\begin{cases}\mu_{i}(\gamma)\cdot\mu_{i+1}(\gamma)=\mu_{i}((\gamma^{v_{i}v_{i+1}})_{v_{i}v_{i+1}})\cdot\mu_{i+1}((\gamma^{v_{i}v_{i+1}})_{v_{i}v_{i+1}})&\\ \mu_{i}(\gamma^{v_{i}v_{i+1}})\cdot\mu_{i+1}(\gamma^{v_{i}v_{i+1}})=\mu_{i}(\gamma_{v_{i}v_{i+1}})\cdot\mu_{i+1}(\gamma_{v_{i}v_{i+1}})&\end{cases}$
      The result then follows.
@@
    If $v$ is not a cut vertex that is a vertex whose removal does not increase the number of components, then it is not too hard to see that there is a natural one-to-one correspondence $\psi$ induced by the merging operations introduced in Section 3 between the set of Euler circuits of $G_{v}$ and the disjoint union of the sets of the Euler circuits of $G_{0}$ and $G_{1}$ In fact, if $\gamma$ is an Euler circuit of $G_{v}$ then $\psi(\gamma)$ is an Euler circuit of $G_{0}$ or $G_{1}$ by choosing the appropriate merging operation at the vertex $v$. Also, if $\beta$ is an Euler circuit of $G_{0}$ or $G_{1}$, then $\psi^{-1}(\beta)$ is an Euler circuit of $G$ obtained by reversing the merging operation at the vertex $v$. Now the chord diagram $C(\psi(\gamma))$ of $\psi(\gamma)$ is obtained from the chord diagram $C(\gamma)$ simply by deleting the $n$-th chord. Thus, if $W(\gamma)=w_{1}(\gamma)\ldots w_{n}(\gamma)$ is the activity word of $\gamma$, then the activity word $W(\psi(\gamma))$ of $\psi(\gamma)$ is $w_{1}(\gamma)\ldots w_{n-1}(\gamma)$. Moreover, it is easy to see that $w_{n}(\gamma)\in\left\{D,d\right\}$ since $v$ is equipped with a positive sign and $C_{n}\neq\emptyset$. According to the proof of Theorem 3.1, the result is independent of the choice of coloring for $G_{v}$. Thus we can choose appropriate coloring of $G_{v}$ in a way that the $n$-th marker in the graph $G_{0}$ is an $A$-marker and the $n$-th marker in the graph $G_{1}$ is an $a$-marker. This implies that $w(\gamma)=qw(\psi(\gamma))$ or $w(\gamma)=q^{-1}w(\psi(\gamma))$, respectively. Thus the result follows directly from the summation over all Euler circuits of $G_{v}$.
@@
    If $v$ is a cut vertex, then one of the two graphs $G_{0}$ or $G_{1}$ consists of two components. Without loss of generality, we can assume that $G_{0}$ is connected and $G_{1}$ consists of two components denoted by $G_{1}^{{}^{\prime}}$ and $G_{1}^{{}^{\prime\prime}}$. Now it is not too hard to see that adding or removing appropriately a new chord that corresponds to the vertex $v$ yields a one-to-one correspondence $\eta$ between the chord diagrams of $G_{v}$ and the chord diagrams of the graph $G_{0}$. Thus, if $W(\gamma)=w_{1}(\gamma)\ldots w_{n}(\gamma)$ is the activity word of $\gamma$, then the activity word $W(\eta(\gamma))$ of $\eta(\gamma)$ is $w_{1}(\gamma)\ldots w_{n-1}(\gamma)$. Moreover, it is easy to see that $w_{n}(\gamma)\in\left\{L,l\right\}$ since $v$ is equipped with a positive sign and $C_{n}(\gamma)=\emptyset$. According to the proof of Theorem 3.1, the result is independent of the choice of coloring for $G_{v}$. Thus we can choose appropriate coloring of $G_{v}$ in a way that the $n$-th marker in the graph $G_{0}$ is an $A$-marker and the $n$-th marker in the graph $G_{1}$ is an $a$-marker. This implies that $X_{G_{v}}(q)=-q^{-3}X_{G_{0}}(q)$. Now we prove that the right-hand side has also the same value. We can see that any chord diagram for the graph $G_{0}$ is a union of two chord diagrams in appropriate way of the graphs $G_{1}^{{}^{\prime}}$ and $G_{1}^{{}^{\prime\prime}}$ and vice-versa. This implies that $X_{G_{1}}(q)=-(q^{2}+q^{-2})X_{G_{0}}(q)$. Now we obtain $qX_{G_{0}}(q)+q^{-1}X_{G_{1}}(q)=qX_{G_{0}}(q)+q^{-1}(-q^{2}-q^{-2})X_{G_{0}}(q)=-q^{-3}X_{G_{0}}(q)$ as requested.

### v2

@@ Abstract
We assign a new polynomial to any vertex-signed checkerboard-colorable 4-valent virtual graph in terms of its Euler circuit expansion. This provides a new combinatorial formulation of the Jones-Kauffman polynomial for checkerboard-colorable virtual links.
footnotetext: 2020 Mathematics Subject Classification. 05C31, 57K14 footnotetext: Key words and phrases. 4-valent graphs, Jones-Kauffman polynomial, Euler circuits, virtual links, checkerboard-colorable.
@@ 1 Introduction
Now we describe the contents of this paper. In Section 2, we briefly introduce and recall the main terms that will be used in this paper. In Section 3, we define the new polynomial for vertex-signed 2-digraphs. Also, we state the main theorem, and we give some properties of this polynomial. In particular, we show that it satisfies a skein relation. At the end of this section, we give an application that shows that the polynomial of the signed-vertex 2-digraph associated to a virtual link diagram allows us to recover its Jones-Kauffman polynomial. The proofs of the primary theorems are addressed in Section 4.
@@ 2 Preliminary
A 2-digraph has a source-target structure iff it is checkerboard-colorable. In fact, if $G$ is a 2-digraph with a source-target structure, then to each oriented edge we associate a normal vector satisfying the right hand rule in the plane as depicted in the left of Figure 3. By coloring the side of each edge containing its normal vector we get a checkerboard coloring of $G$. Conversely, If $G$ is a checkerboard colored 4-valent graph, then, for each edge, we put a normal vector inside its colored side with initial point on that edge as depicted in the right of Figure 3. Then we orient the edge by using the right hand rule.
Throughout the remainder of this paper, we adopt the convention that a checkerboard coloring is associated with a source-target structure as described above.
@@
A 4-valent graph $G$ is said to be checkerboard-embedded in some orientable surface $\Sigma$ if there exists an embedding $e$ of $G$ in $\Sigma$ such that the components of $\Sigma\setminus e(G)$ are 2-cells and there exists a coloring of these components by two colors, say black and white, such that two components of $\Sigma\setminus e(G)$ which are adjacent by an edge of $e(G)$ have always distinct colors. Manturov in [14, 15] showed that a 4-valent virtual graph is checkerboard-colorable iff it is checkerboard-embeddable in some orientable surface.
@@ 2.1 Euler circuits and chord diagrams
In the graph $G$, we let $v_{1},\ldots,v_{n}$ be the set of vertices of such graph and $\gamma$ be an Euler circuit of $G$. Draw a simple circle $C$ in the plane and mark $2n$ evenly spaced points around its circumference. Start traveling $\gamma$ following the orientation of the edges starting from any vertex. Label each of the $2n$ points on the circle $C$ successively with the label of the vertex we encounter. Each vertex label appears twice at exactly two distinct points on the circle. For each vertex $v_{i}$ in $G$, draw a chord inside the circle $C$ connecting the two points labeled $v_{i}$. We ensure that each pair of chords intersect at most once and no more than two chords intersect at the same point. The circle $C$ together with the $n$ chords constitute the chord diagram of $\gamma$, which we denote by $C(\gamma)$. See Figure 4 as an example for an Euler circuit $\gamma$ and its corresponding chord diagram $C(\gamma)$.
@@ 2.3 The Jones-Kauffman polynomial of a virtual link
$L$                       $L_{0}$                      $L_{1}$
Figure 7: The link diagram $L$ at the crossing $c$ and its smoothings $L_{0}$ and $L_{1}$.
For an oriented virtual link diagram $L$, we let $x(L)$ denote the number of negative crossings and $y(L)$ denote the number of positive crossings in $L$ according to the scheme in Figure 8. The writhe of the virtual link diagram $L$ is defined to be the integer $\omega(L)=y(L)-x(L)$.
@@
$f_{L}(q)=(-q)^{-3\omega(L)}\left\langle L\right\rangle\in\mathbb{Z}[q,q^{-1}],$
@@ 3 The main results and an application
We say that a $2$-diagraph is vertex-signed if each vertex is equipped with a sign $+$ or --.
For the rest of this paper, we assume that any $2$-digraph is vertex-signed and checkerboard-colorable unless mentioned otherwise.
Let $G$ be a $2$-diagraph with $n$ vertices $v_{i}$, $1\leq i\leq n$, and $\gamma$ be an Euler circuit of $G$. At each vertex $v_{i}$, we denote the two incoming edges by $e_{i_{1}}$ and $e_{i_{2}}$, and we denote the two outgoing edges by $e_{i_{3}}$ and $e_{i_{4}}$. We assume that at a fixed vertex $v_{i}$ the Euler circuit $\gamma$ of $G$ traverses successively the edges $e_{i_{1}}$ and $e_{i_{3}}$ in the first visit, and then it traverses successively $e_{i_{2}}$ and $e_{i_{4}}$ in the second visit. Merge the edges $e_{i_{1}}$ and $e_{i_{3}}$, i.e., replace them with a single new edge going from the tail of $e_{i_{1}}$ to the head of $e_{i_{3}}$. Also, we merge the edges $e_{i_{2}}$ and $e_{i_{4}}$ using the same technique and we delete the vertex $v_{i}$ to get a new graph with one less vertex. In the vicinity of the deleted vertex $v_{i}$, put a marker connecting the traces of the deleted vertex on each one of the new edges. We label each marker by $A$, $a$, $B$, or $b$ according to the coloring, the sign of the vertex and the order of the edges when traversing the path as explained in Figure 9. Such labeled marker for the vertex $v_{i}$ is denoted by “$i$-th marker”. We iterate this merging process until all the vertices are deleted and we obtain a single immersed circle in the plane with $n$ labeled markers in the vicinities of the deleted vertices. We call such data the $\gamma$-state of the corresponding Euler circuit $\gamma$ of $G$. In Figure 10, we depict the $\gamma$-state of the Euler circuit $\gamma$ given in Figure 4.
Figure 9: The $i$-th marker corresponding to the deleted vertex $v_{i}$ and the rule of its labeling.
@@
where $G_{1},\ldots,G_{m}$ are the connected components of the graph $G$.
Now we state the main theorem of this paper whose proof will be given in the next section.
@@
We enclose this subsection by stating the following theorem whose proof will be given in the next section which implies that the polynomial $X_{G}(q)$ satisfies a skein relation. For this end, we let $G^{v}$ be a checkerboard-colorable vertex-signed $2$-digraph with a fixed vertex $v$ depicted in Figure 11. The digraphs resulting from the two possible merging operations at $v$ described in Section 3 will be denoted by $G_{0}^{v}$ and $G_{1}^{v}$ respectively according to the scheme in Figure 11.
Figure 11: The 2-digraph $G^{v}$ at the fixed vertex $v$ and the corresponding 2-digraphs $G_{0}^{v}$ and $G_{1}^{v}$ respectively obtained by the two merging operations.
@@
    If $v$ is equipped with a positive sign, then
    $X_{G^{v}}(q)=qX_{G_{0}^{v}}(q)+q^{-1}X_{G_{1}^{v}}(q).$
@@
    If $v$ is equipped with a negative sign, then
    $X_{G^{v}}(q)=q^{-1}X_{G_{0}^{v}}(q)+qX_{G_{1}^{v}}(q).$
As an application, we relate the polynomial introduced in this section to the established Jones-Kauffman polynomial of checkerboard-colorable virtual links. We start by recalling a brief background on virtual knot theory focusing on checkeroboard-colorable virtual links. Given a virtual link diagram, one can obtain a 4-valent virtual graph by replacing each classical crossing with a vertex and each virtual crossing by a virtual double point. The obtained graph is called the shadow graph of the given virtual link diagram. If this graph is checkerboard-colorable, then the associated virtual link diagram is checkerboard-colorable. For instance, the virtual link diagram in Figure 5, whose shadow graph is exactly the graph depicted in Figure 1, is checkerboard-colorable. A notable fact about the checkerboard-colorability is that it behaves well under generalized Reidemeister moves since any two equivalent checkerboard-colorable virtual link diagrams are related by a finite sequence of moves which preserve the checkerboard-colorability as shown in [8, Theorem 2.9]. A virtual link is checkerboard-colorable if it admits a checkerboard-colorable diagram. We point out that the notion of checkerboard-colorable virtual links was first introduced by Kamada in [10, 11].
@@
Let $L$ be an oriented checkerboard-colorable virtual link. Let $G$ be the shadow graph of a checkerboard-colorable diagram of $L$, then $f_{L}(q)=(-q)^{-3\omega(L)}X_{G}(q)$, where $f_{L}(q)$ and $\omega(L)$ are the Jones-Kauffman polynomial and the writhe of the oriented virtual link $L$, respectively.
@@
Example 3.5.
We consider the checkerboard colorable knot $K=5.2426$ represented by the diagram on the left of Figure 13. We compute the Jones-Kauffman polynomial of $K$ using our method. The shadow graph of such a knot diagram is depicted on the right of Figure 13.
Figure 13: A diagram of the knot $K=5.2426$ and its shadow graph, respectively.
The following are the chord diagrams of the nine Euler circuits of the shadow graph together with their activity words and weights. We have placed the label of the marker corresponding to each vertex as the exponent of the vertex label.
@@ 4 Proof of the Main Theorems
    The third equivalence is true since either $v_{h}=v_{i}$ or the edges of $H^{v_{i}v_{j}}$ connecting $v_{h}$ to $v_{k}$ are:
@@
      $\begin{cases}\mu_{i}(\gamma)\cdot\mu_{i+1}(\gamma)=\mu_{i}(\gamma_{v_{i}v_{i+1}})\cdot\mu_{i+1}(\gamma_{v_{i}v_{i+1}}),&\\ \mu_{i}(\gamma^{v_{i}v_{i+1}})\cdot\mu_{i+1}(\gamma^{v_{i}v_{i+1}})=\mu_{i}((\gamma^{v_{i}v_{i+1}})_{v_{i}v_{i+1}})\cdot\mu_{i+1}((\gamma^{v_{i}v_{i+1}})_{v_{i}v_{i+1}}).\end{cases}$
      $\quad\text{or}\quad\begin{cases}\mu_{i}(\gamma)\cdot\mu_{i+1}(\gamma)=\mu_{i}((\gamma^{v_{i}v_{i+1}})_{v_{i}v_{i+1}})\cdot\mu_{i+1}((\gamma^{v_{i}v_{i+1}})_{v_{i}v_{i+1}}),&\\ \mu_{i}(\gamma^{v_{i}v_{i+1}})\cdot\mu_{i+1}(\gamma^{v_{i}v_{i+1}})=\mu_{i}(\gamma_{v_{i}v_{i+1}})\cdot\mu_{i+1}(\gamma_{v_{i}v_{i+1}}).\end{cases}$
@@
      $\begin{cases}\mu_{i}(\gamma)\cdot\mu_{i+1}(\gamma)=-\mu_{i}(\gamma^{v_{i}v_{i+1}})\cdot\mu_{i+1}(\gamma^{v_{i}v_{i+1}}),&\\ \mu_{i}(\gamma_{v_{i}v_{i+1}})\cdot\mu_{i+1}(\gamma_{v_{i}v_{i+1}})=-\mu_{i}((\gamma^{v_{i}v_{i+1}})_{v_{i}v_{i+1}})\cdot\mu_{i+1}((\gamma^{v_{i}v_{i+1}})_{v_{i}v_{i+1}}).\end{cases}$
@@
      $\begin{cases}\mu_{i}(\gamma)\cdot\mu_{i+1}(\gamma)=\mu_{i}((\gamma^{v_{i}v_{i+1}})_{v_{i}v_{i+1}})\cdot\mu_{i+1}((\gamma^{v_{i}v_{i+1}})_{v_{i}v_{i+1}}),&\\ \mu_{i}(\gamma^{v_{i}v_{i+1}})\cdot\mu_{i+1}(\gamma^{v_{i}v_{i+1}})=\mu_{i}(\gamma_{v_{i}v_{i+1}})\cdot\mu_{i+1}(\gamma_{v_{i}v_{i+1}}).\end{cases}$
      Hence the result then follows.
@@
    If $v$ is not a cut vertex that is a vertex whose removal does not increase the number of components, then it is not too hard to see that there is a natural one-to-one correspondence $\psi$ induced by the merging operations introduced in Section 3 between the set of Euler circuits of $G^{v}$ and the disjoint union of the sets of the Euler circuits of $G_{0}^{v}$ and $G_{1}^{v}$ In fact, if $\gamma$ is an Euler circuit of $G^{v}$ then $\psi(\gamma)$ is an Euler circuit of $G_{0}^{v}$ or $G_{1}^{v}$ by choosing the appropriate merging operation at the vertex $v$. Also, if $\beta$ is an Euler circuit of $G_{0}^{v}$ or $G_{1}^{v}$, then $\psi^{-1}(\beta)$ is an Euler circuit of $G$ obtained by reversing the merging operation at the vertex $v$. Now the chord diagram $C(\psi(\gamma))$ of $\psi(\gamma)$ is obtained from the chord diagram $C(\gamma)$ simply by deleting the $n$-th chord. Thus, if $W(\gamma)=w_{1}(\gamma)\ldots w_{n}(\gamma)$ is the activity word of $\gamma$, then the activity word $W(\psi(\gamma))$ of $\psi(\gamma)$ is $w_{1}(\gamma)\ldots w_{n-1}(\gamma)$. Moreover, it is easy to see that $w_{n}(\gamma)\in\left\{D,d\right\}$ since $v$ is equipped with a positive sign and $C_{n}\neq\emptyset$. According to the proof of Theorem 3.1, the result is independent of the choice of coloring for $G^{v}$. Thus we can choose appropriate coloring of $G^{v}$ in a way that the $n$-th marker in the graph $G_{0}^{v}$ is an $A$-marker and the $n$-th marker in the graph $G_{1}^{v}$ is an $a$-marker. This implies that $w(\gamma)=qw(\psi(\gamma))$ or $w(\gamma)=q^{-1}w(\psi(\gamma))$, respectively. Thus the result follows directly from the summation over all Euler circuits of $G^{v}$.
@@
    If $v$ is a cut vertex, then one of the two graphs $G_{0}^{v}$ or $G_{1}^{v}$ consists of two components. Without loss of generality, we can assume that $G_{0}^{v}$ is connected and $G_{1}^{v}$ consists of two components denoted by $(G_{1}^{v})^{\prime}$ and $(G_{1}^{v})^{\prime\prime}$. Now it is not too hard to see that adding or removing appropriately a new chord that corresponds to the vertex $v$ yields a one-to-one correspondence $\eta$ between the chord diagrams of $G^{v}$ and the chord diagrams of the graph $G_{0}^{v}$. Thus, if $W(\gamma)=w_{1}(\gamma)\ldots w_{n}(\gamma)$ is the activity word of $\gamma$, then the activity word $W(\eta(\gamma))$ of $\eta(\gamma)$ is $w_{1}(\gamma)\ldots w_{n-1}(\gamma)$. Moreover, it is easy to see that $w_{n}(\gamma)\in\left\{L,l\right\}$ since $v$ is equipped with a positive sign and $C_{n}(\gamma)=\emptyset$. According to the proof of Theorem 3.1, the result is independent of the choice of coloring for $G^{v}$. Thus we can choose appropriate coloring of $G^{v}$ in a way that the $n$-th marker in the graph $G_{0}^{v}$ is an $A$-marker and the $n$-th marker in the graph $G_{1}^{v}$ is an $a$-marker. This implies that $X_{G^{v}}(q)=-q^{-3}X_{G_{0}^{v}}(q)$. Now we prove that the right-hand side has also the same value. We can see that any chord diagram for the graph $G_{0}^{v}$ is a union of two chord diagrams in appropriate way of the graphs $(G_{1}^{v})^{\prime}$ and $(G_{1}^{v})^{\prime\prime}$ and vice-versa. This implies that $X_{G_{1}^{v}}(q)=-(q^{2}+q^{-2})X_{G_{0}^{v}}(q)$. Now we obtain $qX_{G_{0}^{v}}(q)+q^{-1}X_{G_{1}^{v}}(q)=qX_{G_{0}^{v}}(q)+q^{-1}(-q^{2}-q^{-2})X_{G_{0}^{v}}(q)=-q^{-3}X_{G_{0}^{v}}(q)$ as requested.